Circular No. 7814

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2001 OG_{108} (LONEOS) An apparently asteroidal object reported by LONEOS (discovery ob-

servation below) was announced on MPEC 2001-P40 last Aug. 13. Numerous observers could not detect any cometary activity, despite the cometlike orbit. Now approaching perihelion, the object has finally shown cometary activity, detected by several observers as indicated by the following reported total magnitudes and coma/tail data from CCD images: Jan. 11.44 UT, 16.1, 0′.4 coma, 0′.5 tail in p.a. 230^{o} (A. Nakamura, Kuma, Ehime, Japan, 0.60-m reflector); 22.39, 14.7, diffuse (S. Wakuda, Shizuoka, Japan, 0.25-m reflector); Feb. 1.40, 13.5, 0′.9 coma, broad tail toward west with the longest segment being 0′.9 long in p.a. 220^{o} (Nakamura); 1.41, 12.8, fan-shaped coma spanning p.a. 200^{o}–345^{o} (T. Oribe, Saji, Tottori, Japan, 1.03-m reflector). Recent astrometry and the orbital elements below appear on MPEC 2002-C04.

2001 UT  α_{2000}  δ_{2000}  m_{2}
July 28.39388  1^{h}46^{m}17.^{s}02  +10^{o}53′25′′.4  18.7

Epoch = 2002 Mar. 27.0 TT

T = 2002 Mar. 15.20618 TT  ω = 116.^{o}41991
e = 0.9252940  Ω = 10.55525 }2000.0
q = 0.9940480 AU  i = 80.24516

a = 13.3061362 AU n^{o} = 0.02030608 P = 48.538 years

2002 TT α_{2000}  δ_{2000}  Δ  r  ϵ  β  m_{1}
Jan. 26 21^{h}30.^{m}13 +28^{o}28.^{′}0 1.663 1.264 49.^{o}3 36.^{o}2 13.1
Feb. 5 21 30.99 +31 14.3 1.588 1.174 47.5 38.2 12.7
15 21 32.24 +34 41.8 1.481 1.097 47.8 41.8 12.3
25 21 33.72 +39 02.8 1.342 1.039 50.1 47.0 11.8
Mar. 7 21 35.61 +44 40.0 1.176 1.003 54.3 53.5 11.4
17 21 38.74 +52 20.8 0.990 0.994 60.1 60.2 11.0
27 21 46.01 +63 39.6 0.801 1.013 67.5 65.5 10.6

COMET C/2000 WM_{1} (LINEAR) Total naked-eye magnitude estimates of this comet (cf. IAUC 7809):

Jan. 29.78 UT, 3.0 (M. Mattiazzo, Wallaroo, S. Australia); 30.85, 2.8 (A. Pearce, Nedlands, W. Australia); 31.31, 3.2 (J. G. de S. Aguiar, Campinas, Brazil); Feb. 2.31, 3.5 (A. Amorim, Florianopolis, Brazil; moonlight).

2002 February 2  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT  Daniel W. E. Green